Circular No. 8162

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

THE EDGAR WILSON AWARD The Smithsonian Astrophysical Observatory announces that the 2003

Edgar Wilson Award for the discovery of comets (cf. IAUC 6936, 7926) is being divided among the following five individuals: Sebastian Florian Hönig, Dossenheim, Germany, for C/2002 O4; Tetuo Kudo, Kikuchi, Kumamoto, Japan, and Shigehisa Fujikawa, Mitoyo, Kagawa, Japan, for C/2002 X5; Charles Wilson Juels, Fountain Hills, AZ, U.S.A., and Paulo Renato Centeno Holvorcem, Campinas, Brazil, for C/2002 Y1.

SUPERNOVA 2003gk IN NGC 7460 Further to IAUC 8161, J. Graham and W. Li report the LOTOSS

discovery of an apparent supernova on unfiltered KAIT images taken on July 1.5 (mag ∼ 17.1) and July 2.5 UT (mag ∼ 17.2). SN 2003gk is located at α = 23^{h}01^{m}42^{s}.99, δ = +2^{o}16′08′′.7 (equinox 2000.0), which is 1′′.6 east and 19′′.2 north of the nucleus of NGC 7460. A KAIT image taken on 2002 Dec. 3.2 showed nothing at this position (limiting mag ∼ 19.0).

MARS J. Beish, Lake Placid, FL (0.41-m reflector, visual); and D. C. Parker,

Coral Gables, FL (0.41-m reflector, CCD), report that a dust cloud has formed and brightened during July 1 and 2 in the Iapygia-Ionium-Hellas region centered at longitude ≈ 294^{o} west and latitude ≈ 25^{o} south.

η CARINAE

Z. Abraham and T. P. Dominici, São Paulo University and São Paulo State Research Foundation, report that radio measurements at 7 mm, made at the Itapetinga Radio Observatory, show that the flux density decreased from 3 Jy on May 22 to 1.5 Jy on June 20, remaining constant since then. A. Damineli, N. Leister, and R. Levenhagen report that spectroscopy, obtained at the Pico dos Dias Observatory, shows the complete disappearance of the [Ne III] 386.8-nm spectral line on June 23. Both results indicate that η Car has entered the deepest phase of the 2003.5 event, as predicted by Damineli et al. (2000, Ap.J. 528, L101).

SUPERNOVA 2003gd IN M74 CCD photometry by M. Lehky and K. Hornoch, Brno, Czech Rep.:

June 29.037 UT, V = 14.05 ± 0.11; 29.046, R = 13.68; 29.054, I = 13.27.

2003 July 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green